# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v2

@@ Abstract
We study the problem of efficiently computing the derivative of the fixed-point of a parametric nondifferentiable contraction map. This problem has wide applications in machine learning, including hyperparameter optimization, meta-learning and data poisoning attacks. We analyze two popular approaches: iterative differentiation (ITD) and approximate implicit differentiation (AID). A key challenge behind the nonsmooth setting is that the chain rule does not hold anymore. Building upon the recent work by Bolte et al. (2022), who proved linear convergence of nondifferentiable ITD, we provide an improved linear rate for ITD and a slightly better rate for AID, both in the deterministic case. We further introduce NSID, a new stochastic method to compute the implicit derivative when the fixed point is defined as the composition of an outer map and an inner map which is accessible only through a stochastic unbiased estimator. We establish rates for the convergence of NSID, encompassing the best available rates in the smooth setting. We present illustrative experiments confirming our analysis.
@@ 1 Introduction
when $\Phi$ is not differentiable, but only differentiable almost everywhere. We address both the case that $\Phi$ can be explicitly evaluated, and the case that $\Phi$ has the composite form
@@
Recently Bolte & Pauwels (2021) introduced the notion of conservative derivatives as an effective tool to rigorously address automatic differentiation of neural networks with nondifferentiable activations (e.g., ReLU). Moreover, if $\Phi(\cdot,\lambda)$ is a contraction and under the general assumption that $\Phi$ is piecewise Lipschitz smooth with finite pieces, Bolte et al. (2022) provide an asymptotic linear convergence rate for deterministic ITD. [^1] Therein, referred to as piggyback automatic differentiation. However, we are not aware of any result of such type for the AID method and for the stochastic setting of problem (2), even when $G(v,\lambda)=v$. In particular the compositional structure (2) allows us to cover e.g., proximal stochastic gradient methods, which are a common and practical example of nonsmooth optimization algorithms, but it adds additional challenges since we do not have access to an unbiased estimator of $\Phi$ as for the smooth stochastic case studied in (Grazzi et al., 2021, 2023).
@@
  • We prove non-asymptotic linear convergence rates for deterministic ITD and AID which, from one hand extend the results for the case where $\Phi$ is Lipschitz smooth given in (Grazzi et al., 2020), and on the other end, improve the result in (Bolte et al., 2022) for nonsmooth ITD. The given bounds indicate that AID converges faster than ITD, which we verify empirically. We also identify cases in which this difference in performance in favor of AID might be large due to nondifferentiable regions.
  • We propose the first stochastic AID approach with proven convergence rates, which we name nonsmooth stochastic implicit differentiation (NSID). Notably, we prove that NSID can converge to a true conservative Jacobian-vector product with rate $O(1/k)$, where $k$ is the number of samples, provided that the fixed-point problem is solved with rate $O(1/k)$.
@@
When $\Phi$ is differentiable and under some regularity assumptions, approximation guarantees have been established for AID and ITD approaches in the deterministic setting (Pedregosa, 2016; Grazzi et al., 2020), and for AID approaches in the special case of the stochastic setting (2) where $G(v,\lambda)=v$ and we have access only to $\hat{T}$ (Grazzi et al., 2021, 2023). Furthermore, several works established convergence rates and, in the stochastic setting, sample complexity results for bilevel optimization algorithms relying on AID and ITD approaches, see e.g., (Ghadimi & Wang, 2018; Ji et al., 2021; Arbel & Mairal, 2021; Chen et al., 2021).
@@ 2 Preliminaries
$\mathcal{A}(\mathcal{X}){:=}\mathcal{A}\,\begin{bmatrix}\mathcal{X}\\ I_{p_{2}}\end{bmatrix}{:=}\big{\{}A_{1}X+A_{2}\,\big{|}\,[A_{1}A_{2}]\in\mathcal{A},X\in\mathcal{X}\big{\}},$ (3)
@@
For any integer $r\geq 1$ we set $[r]=\{1,\dots,r\}$. If $F\colon\mathbb{R}^{p_{1}+p_{2}}\to\mathbb{R}^{d}$ is differentiable, we denote by $F^{\prime}(x)\in\mathbb{R}^{d\times(p_{1}+p_{2})}$ the derivative of $F$ (its Jacobian) at $x$ and by $\partial_{1}F(x)\in\mathbb{R}^{d\times p_{1}}$ and $\partial_{2}F(x)\in\mathbb{R}^{d\times p_{2}}$ the partial derivatives of $F$ with respect to the first and second block of variables respectively. For a random vector $\xi\in\mathbb{R}^{d}$, we denote with $\mathbb{E}[\xi]$ its expectation and with $\operatorname{Var}[\xi]=\mathbb{E}\lVert\xi-\mathbb{E}[\xi]\rVert^{2}$ its variance. In our assumptions we will consider the class of so called definable functions, which includes the large majority of functions used for machine learning applications (see Appendix A).
@@ 2.1 Conservative Derivatives
Let $U\subset\mathbb{R}^{p}$ be an open set and $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a locally Lipschitz continuous mapping. We say that a set-valued mapping $D_{F}\colon U\rightrightarrows\mathbb{R}^{d\times p}$ is a conservative derivative (CD) of $F$, if $D_{F}$ has closed graph, nonempty compact values, and for every absolutely continuous curve $\gamma\colon[0,1]\to U\subset\mathbb{R}^{p}$ we have that, for almost every $t\in[0,1]$
@@
Conservative derivatives are extensively analyzed in (Bolte & Pauwels, 2021). Some key properties are that: (1) they are almost everywhere single-valued and equal to classical derivatives; (2) for path differentiable functions, the Clarke subgradient is the minimal conservative derivative up to a convex envelope; (3) chain rule holds for conservative derivatives; (4) locally Lipschitz definable mappings admit conservative derivatives. We also point out that – as it is usual for generalized derivatives – conservative derivatives are unique only up to a set of Lebesgue measure zero. This accounts for the fact that there are multiple ways to express a path differentiable function as a composition of others but applying the chain rule produces always valid CDs.
@@
Note that $\lVert\mathcal{A}\rVert_{\sup}=\operatorname{\it{e}}(\mathcal{A},\{0\})$. The excess satisfies several properties similar to the ones of a distance, even though it is not symmetric (see Lemma B.1). Similarly to (Scholtes, 2012) we give the following concept of piecewise continuity and smoothness (which is slightly more general than that given in (Bolte & Pauwels, 2021)).
@@ 3 Differentiating a Parametric Fixed Point
where $f_{\lambda}=\mathbb{E}\hat{f}_{\lambda}(\cdot,\xi)$ is convex $L$-smooth, while $g_{\lambda}$ is convex lower semicontinuous extended-real valued. This can be cast into (2) by setting $\eta\in\left[0,2/L\right[$, $\hat{T}_{\xi}(w,\lambda)=w-\eta\hat{\nabla}f_{\lambda}(w,\xi)$ and $G(w,\lambda)=\operatorname{Prox}_{\eta g_{\lambda}}(w)$ with $\operatorname{Prox}_{h}(x)=\operatorname*{arg\,min}_{y}(h(x)+(1/2)\lVert x-y\rVert^{2})$ being the proximity operator of $h$. Several machine learning problems can be written in form (8), e.g., LASSO, elastic net, (dual) SVM, where $g_{\lambda}$ is not smooth.
@@
Our theoretical analysis requires only that $\Phi$ is definable piecewise smooth and that the inequality in Lemma 3.2 holds for some conservative derivatives of $\Phi$, even if it is not computed according to (10). One such situation occurs for instance when $\Phi$ has the structure of a finite sum, that is, $\Phi=\sum_{i=1}^{n}\Phi^{(i)}$, where each $\Phi^{(i)}$ satisfies Assumption 3.1(i) with corresponding conservative derivative $D_{\Phi^{(i)}}^{s}$. Then, it is clear that $\Phi$ is still definable and piecewise Lipschitz smooth. Moreover, using the properties of conservative derivatives (see Corollary 4 in (Bolte & Pauwels, 2020)), $D_{\Phi}=\sum_{i=1}^{n}D_{\Phi^{(i)}}^{s}$ is a conservative derivatives of $\Phi$. Thus, using the property of the excess (see Lemma B.1(ii)) it directly follows that the inequality in Lemma 3.2, and hence our theory, still holds for such $\Phi$.
@@ 4 Deterministic Iterative and Approximate Implicit Differentiation
In the Lipschitz smooth case Grazzi et al. (2020) proved non-asymptotic linear convergence rates for both methods, revealing that AID-FP is slightly faster than ITD. We now extend this analysis to nonsmooth ITD and AID-FP, focusing on the convergence of the set-valued iterations in (14) and (15). Thanks to Lemma 3.2 and the properties of the excess, the proof (in Appendix C) can proceed similarly to that given in (Grazzi et al., 2020) for the smooth case.
@@ 5 Nonsmooth Stochastic Implicit Differentiation
    $T$ and $G$ satisfy Assumption 3.1(i) individually, with constant $L_{T}$ and $L_{G}$ respectively. Let also $T^{\prime}$ and $G^{\prime}$ be selections of the conservative derivatives $D_{T}$ and $D_{G}$ respectively. Also, $\Phi(u,\lambda)=G(T(u,\lambda),\lambda)$.
@@
5.1 ensures that $D_{\Phi}$ obtained via the chain rule for conservative derivatives in (Bolte & Pauwels, 2021) (see Appendix D) is a conservative derivative of $\Phi$ and that $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$. Thus, $w(\lambda)$ is well defined and it has conservative derivatives $D_{w}^{\operatorname{imp}}$ and $D_{w}^{\operatorname{fix}}$. 5.2 is a nonsmooth generalization of the corresponding one in (Grazzi et al., 2021, 2023). Finally, recalling (11), if $\partial_{2}\Phi(u,\lambda)=\partial_{1}G(T(u,\lambda),\lambda)\partial_{2}T(u,\lambda)+\partial_{2}G(T(u,\lambda),\lambda)$ then
@@
with $\bar{T}^{\prime}(u,\lambda){:=}[\partial_{1}\bar{T}(u,\lambda),\partial_{1}\bar{T}(u,\lambda)]{=}\frac{1}{J}\sum_{j=1}^{J}\hat{T}^{\prime}_{\xi_{j}^{(1)}}(u,\lambda)$, which thanks to the chain rule is an element of a partial conservative derivative of $\bar{\Phi}$ (see also Appendix D).
@@
If $G(\cdot,\lambda)$ is the identity and $T$ is smooth, NSID reduces to the same procedure given in (Grazzi et al., 2023), which also provide the bound $O\left(\sigma_{\lambda}(k)+\kappa^{2}J^{-1}+\kappa^{4}\rho_{\lambda}(t)\right)$ in Theorem 7. Compared to the bound given in Theorem 5.5, we note that the only difference is in the constant in front of the term $J^{-1}$, which we believe may be related to the term $G$. Indeed handling a general $G$ provides an additional challenge since we do not have access anymore to an unbiased estimator of $\Phi$. However, we could overcome this issue by using different samples sequences for the two factors occuring in $\hat{\Psi}$. Incidentally, one of those can be the one used to compute a mini-batch estimator of $\partial_{2}\Phi$. Ultimately, this does not call for any additional samples compared to the smooth version, but it could worsen some constants in the bound.
@@ 6 Application to Bilevel Optimization
is a conservative derivatives for $f$. We also let $f_{t}(\lambda):=E(w_{t}(\lambda),\lambda)$, where $w_{t}(\lambda)$ is an approximate solution for the fixed point problem.
@@
Despite these encouraging results and the fact that in the smooth case several works provide convergence rates to a stationary point of the gradient of $f$ (Ji et al., 2021; Arbel & Mairal, 2021; Grazzi et al., 2023), proving such type of results or even asymptotic convergence (without rates) in our nonsmooth case is more challenging and we leave it for future work. One crucial issue is that in the analysis, the constant defined in Lemma 2.5, which we use in place of that of Lipschitz smoothness, cannot be properly controlled on the whole $\Lambda$ as required in the smooth case: it becomes arbitrarily large when $(w(\lambda),\lambda)$ approaches nondifferentiable regions of $\Phi$.

### v3

@@ Abstract
We study the problem of efficiently computing the derivative of the fixed-point of a parametric nondifferentiable contraction map. This problem has wide applications in machine learning, including hyperparameter optimization, meta-learning and data poisoning attacks. We analyze two popular approaches: iterative differentiation (ITD) and approximate implicit differentiation (AID). A key challenge behind the nonsmooth setting is that the chain rule does not hold anymore. We build upon the work by Bolte et al. (2022), who prove linear convergence of nonsmooth ITD under a piecewise Lipschitz smooth assumption. In the deterministic case, we provide a linear rate for AID and an improved linear rate for ITD which closely match the ones for the smooth setting. We also introduce NSID, a new stochastic method to compute the implicit derivative when the contraction map is defined as the composition of an outer map and an inner map which is accessible only through a stochastic unbiased estimator. We establish rates for the convergence of NSID, encompassing the best available rates in the smooth setting. We present illustrative experiments confirming our analysis.
@@ 1 Introduction
when $\Phi$ is not differentiable, but only piecewise differentiable. We address both the case that $\Phi$ can be explicitly evaluated, and the case that $\Phi$ has the composite form
@@
Recently Bolte & Pauwels (2021) introduced the notion of conservative derivatives as an effective tool to rigorously address automatic differentiation of neural networks with nondifferentiable activations (e.g., ReLU). Moreover, if $\Phi(\cdot,\lambda)$ is a contraction and under the general assumption that $\Phi$ is piecewise Lipschitz smooth with finite pieces, Bolte et al. (2022) provide an asymptotic linear convergence rate for deterministic ITD. [^1] Therein, referred to as piggyback automatic differentiation. However, such rate is worse than that of the smooth case and we are not aware of any result of such type for the AID method and for the stochastic setting of problem (2), even when $G(v,\lambda)=v$. In particular the compositional structure (2) allows us to cover e.g., proximal stochastic gradient methods, which are a common and practical example of nonsmooth optimization algorithms, but it adds additional challenges since we do not have access to an unbiased estimator of $\Phi$ as for the smooth stochastic case studied in (Grazzi et al., 2021, 2023).
@@
  • We prove non-asymptotic linear convergence rates for deterministic ITD and AID which, from one hand extend the rates for the case where $\Phi$ is Lipschitz smooth given in (Grazzi et al., 2020), which are fully recovered as a special case, and on the other end, improve the result in (Bolte et al., 2022) for nonsmooth ITD. The given bounds indicate that AID converges faster than ITD, which we verify empirically. We also identify cases in which this difference in performance in favor of AID might be large due to nondifferentiable regions.
  • We propose the first nonsmooth stochastic AID approach with proven convergence rates, which we name nonsmooth stochastic implicit differentiation (NSID). Notably, we prove that NSID can converge to a true conservative Jacobian-vector product with rate $O(1/k)$, where $k$ is the number of samples, provided that the fixed-point problem is solved with rate $O(1/k)$.
@@
When $\Phi$ is differentiable and under some regularity assumptions, approximation guarantees have been established for AID and ITD approaches in the deterministic setting (Pedregosa, 2016; Grazzi et al., 2020), and for AID in the special case of the stochastic setting (2) where $G(v,\lambda)=v$ (Grazzi et al., 2021, 2023). Furthermore, several works established convergence rates and, in the stochastic setting, sample complexity results for bilevel optimization algorithms relying on AID and ITD approaches, see e.g., (Ghadimi & Wang, 2018; Ji et al., 2021; Arbel & Mairal, 2021; Chen et al., 2021).
@@ 2 Preliminaries
$\mathcal{A}(\mathcal{X}){:=}\mathcal{A}\,\begin{bmatrix}\mathcal{X}\\ I_{p_{2}}\end{bmatrix}{:=}\big{\{}A_{1}X+A_{2}\,\big{|}\,[A_{1},A_{2}]\in\mathcal{A},X\in\mathcal{X}\big{\}},$ (3)
@@
For any integer $r\geq 1$ we set $[r]=\{1,\dots,r\}$. If $F\colon\mathbb{R}^{p_{1}+p_{2}}\to\mathbb{R}^{d}$ is differentiable, we denote by $F^{\prime}(x)\in\mathbb{R}^{d\times(p_{1}+p_{2})}$ the derivative of $F$ (its Jacobian) at $x$ and by $\partial_{1}F(x)\in\mathbb{R}^{d\times p_{1}}$ and $\partial_{2}F(x)\in\mathbb{R}^{d\times p_{2}}$ the partial derivatives of $F$ with respect to the first and second block of variables respectively. Let $F:U\subset\mathbb{R}^{d}\to\mathbb{R}^{p}$, we say that $F$ is smooth (nonsmooth) if it is differentiable (not differentiable) and Lipschitz smooth with constant $L$ or $L$-smooth if it is smooth and $\exists L>0$ such that $\forall x,y\in U$ $\lVert F^{\prime}(x)-F^{\prime}(y)\rVert\leq L\lVert x-y\rVert$. For a random vector $\xi\in\mathbb{R}^{d}$, we denote with $\mathbb{E}[\xi]$ its expectation and with $\operatorname{Var}[\xi]=\mathbb{E}\lVert\xi-\mathbb{E}[\xi]\rVert^{2}$ its variance. In our assumptions we will consider the class of the so called definable functions, which includes the large majority of functions used for machine learning applications (see Appendix A).
@@ 2.1 Conservative Derivatives
Let $U\subset\mathbb{R}^{p}$ be an open set and $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a locally Lipschitz continuous mapping. We say that a set-valued mapping $D_{F}\colon U\rightrightarrows\mathbb{R}^{d\times p}$ is a conservative derivative of $F$, if $D_{F}$ has closed graph, nonempty compact values, and for every absolutely continuous curve $\gamma\colon[0,1]\to U\subset\mathbb{R}^{p}$ we have that, for almost every $t\in[0,1]$
@@
Conservative derivatives are extensively analyzed in (Bolte & Pauwels, 2021). Some key properties are that: (1) they are almost everywhere single-valued and equal to classical derivatives; (2) for path differentiable functions, the Clarke subgradient is the minimal conservative derivative up to a convex envelope; (3) chain rule holds for conservative derivatives; (4) locally Lipschitz definable mappings admit conservative derivatives. We also point out that – as it is usual for generalized derivatives – conservative derivatives are unique only up to a set of Lebesgue measure zero. This accounts for the fact that there are multiple ways to express a path differentiable function as a composition of others but applying the chain rule produces conservative derivatives that can differ but are always valid.
@@
Note that $\operatorname{\it{e}}(\mathcal{A},\mathcal{B})=0\implies\mathcal{A}\subset\mathcal{B}$ and $\lVert\mathcal{A}\rVert_{\sup}=\operatorname{\it{e}}(\mathcal{A},\{0\})$. The excess satisfies several properties similar to the ones of a distance, even though it is not symmetric (see Lemma B.1). Similarly to (Scholtes, 2012) we give the following concept of piecewise continuity and smoothness (which is slightly more general than that given in (Bolte & Pauwels, 2021)).
@@ 3 Differentiating a Parametric Fixed Point
where $f_{\lambda}=\mathbb{E}\hat{f}_{\lambda}(\cdot,\xi)$ is convex $L$-smooth, while $g_{\lambda}$ is convex lower semicontinuous extended-real valued. This can be cast into (2) by setting $\eta\in\left[0,2/L\right[$, $\hat{T}_{\xi}(w,\lambda)=w-\eta\hat{\nabla}f_{\lambda}(w,\xi)$ and $G(w,\lambda)=\operatorname{Prox}_{\eta g_{\lambda}}(w)$ with $\operatorname{Prox}_{h}(x)=\operatorname*{arg\,min}_{y}(h(x)+(1/2)\lVert x-y\rVert^{2})$ being the proximity operator of $h$. Several machine learning problems can be written in form (8) where $g_{\lambda}$ is nonsmooth, e.g., LASSO, elastic net, (dual) SVM.
@@
Our theoretical analysis requires only that $\Phi$ is definable piecewise smooth and that the inequality in Lemma 3.2 holds for some conservative derivatives of $\Phi$, even if it is not computed according to (10). One such situation occurs for instance when $\Phi$ has the structure of a finite sum, that is, $\Phi=\sum_{i=1}^{n}\Phi^{(i)}$, where each $\Phi^{(i)}$ satisfies Assumption 3.1(i) with corresponding conservative derivative $D_{\Phi^{(i)}}^{s}$. Then, it is clear that $\Phi$ is still definable and piecewise Lipschitz smooth. Moreover, using the properties of conservative derivatives (see Corollary 4 in (Bolte & Pauwels, 2020)), $D_{\Phi}=\sum_{i=1}^{n}D_{\Phi^{(i)}}^{s}$ is a conservative derivative of $\Phi$. Thus, using the property of the excess (see Lemma B.1(ii)) it directly follows that the inequality in Lemma 3.2, and hence our theory, still holds for such $\Phi$.
@@ 4 Deterministic Iterative and Approximate Implicit Differentiation
In the Lipschitz smooth case Grazzi et al. (2020) proved non-asymptotic linear convergence rates for both methods, revealing that AID-FP is slightly faster than ITD. We now extend this analysis to nonsmooth ITD and AID-FP, focusing on the convergence of the set-valued iterations in (14) and (15). Thanks to Lemma 3.2 and the properties of the excess, the proof (in Appendix C) can proceed similarly to the one for the smooth case.
@@ 5 Nonsmooth Stochastic Implicit Differentiation
    $T$ and $G$ satisfy Assumption 3.1(i) individually, with constant $L_{T}$ and $L_{G}$ respectively. Let $T^{\prime}$ and $G^{\prime}$ be selections of the conservative derivatives $D_{T}$ and $D_{G}$ respectively. Also, $\Phi(u,\lambda)=G(T(u,\lambda),\lambda)$.
@@
5.1 ensures that $D_{\Phi}$ obtained via the chain rule for conservative derivatives in (Bolte & Pauwels, 2021) (see Appendix D) is a conservative derivative of $\Phi$ and that $\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\leq q<1$. Thus, $w(\lambda)$ is well defined and it has conservative derivatives $D_{w}^{\operatorname{imp}}$ and $D_{w}^{\operatorname{fix}}$. 5.2 is a nonsmooth generalization of the corresponding one in (Grazzi et al., 2021, 2023). Finally, recalling (11), if we set $\partial_{2}\Phi(u,\lambda)=\partial_{1}G(T(u,\lambda),\lambda)\partial_{2}T(u,\lambda)+\partial_{2}G(T(u,\lambda),\lambda)$ then
@@
with $\bar{T}^{\prime}(u,\lambda){:=}[\partial_{1}\bar{T}(u,\lambda),\partial_{2}\bar{T}(u,\lambda)]{=}\frac{1}{J}\sum_{j=1}^{J}\hat{T}^{\prime}_{\xi_{j}^{(1)}}(u,\lambda)$, which thanks to the chain rule is an element of a partial conservative derivative of $\bar{\Phi}$ (see also Appendix D).
@@
If $G(\cdot,\lambda)$ is the identity and $T$ is smooth, NSID reduces to the same procedure given in (Grazzi et al., 2023), which also provide the bound $O\left(\sigma_{\lambda}(k)+\kappa^{2}J^{-1}+\kappa^{4}\rho_{\lambda}(t)\right)$ in Theorem 7. Compared to the bound given in Theorem 5.5, we note that the only difference is in the constant in front of the term $J^{-1}$, which we believe may be related to the term $G$. Indeed handling a general $G$ provides an additional challenge since we do not have access anymore to an unbiased estimator of $\Phi$. However, we could overcome this issue by using different samples sequences for the two factors occuring in $\hat{\Psi}$. Incidentally, one of those sequences can be the one used to compute a mini-batch estimator of $\partial_{2}\Phi$. Ultimately, this does not call for any additional samples compared to the smooth version, but it could worsen some constants in the bound.
@@ 6 Application to Bilevel Optimization
is a conservative derivative for $f$. We also let $f_{t}(\lambda):=E(w_{t}(\lambda),\lambda)$, where $w_{t}(\lambda)$ is an approximate solution for the fixed point problem.
@@
Despite these encouraging results and the fact that in the smooth case several works provide convergence rates to a stationary point of the gradient of $f$ (Ji et al., 2021; Arbel & Mairal, 2021; Grazzi et al., 2023, and others), proving such type of results or even asymptotic convergence (without rates) in our nonsmooth case is more challenging and we leave it for future work. One crucial issue is that in the analysis, the constant defined in Lemma 2.5, which we use in place of that of Lipschitz smoothness, cannot be properly controlled on the whole $\Lambda$ as required in the smooth case: it becomes arbitrarily large when $(w(\lambda),\lambda)$ approaches nondifferentiable regions of $\Phi$.